OpenStudy (anonymous):

antiderivative of (2+x^2)/(1+x^2)

zepdrix (zepdrix):

\[\Large \int\limits \frac{2+x^2}{1+x^2}\;dx \quad=\quad \frac{1+\color{#DD4747}{1+x^2}}{\color{#DD4747}{1+x^2}}\;dx\]

zepdrix (zepdrix):

\[\Large =\quad\int\limits \frac{1}{1+x^2}+\color{#DD4747}{1}\quad dx\]

zepdrix (zepdrix):

Understand how I simplified the fraction? :o

OpenStudy (anonymous):

yes

zepdrix (zepdrix):

Hmm have you learned about this integral yet?\[\Large \int\limits \frac{1}{1+x^2}dx\]

OpenStudy (anonymous):

arctanx

zepdrix (zepdrix):

good good good.

OpenStudy (anonymous):

so would it be arctanx +x +c ?

zepdrix (zepdrix):

yes good job